OpenStudy (anonymous):

It says magnetic flux density is 0.9 w/msquared. What doz it mean?

OpenStudy (festinger):

Magnetic flux is the amount of magnetic field lines that pass perpendicular a surface of an area. Magnetic flux is the average amount of magnetic field lines that pass a surface per unit area. Magnetic flux is measured in T m^2 or Wb. Magnetic flux density is Wb/m^2

OpenStudy (anonymous):

A straight conducting rod of length 0.4m is moving perpendicular to a magnetic flux density of 0.9 w/m2 with a speed of 7 m/s. The emf developed in the rod is

OpenStudy (festinger):

\[|E|=vBL=(7\frac{m}{s})(0.9\frac{Wb}{m^{2}})(0.4m)=2.52V\]

OpenStudy (anonymous):

thnx

OpenStudy (ash2326):

@Festinger Please explain, how did you work through the problem and got the answer. Just using the formula and giving out the answer won't help the user.

OpenStudy (festinger):

It is an experimental fact that \[E=\frac{-d\Phi_{b}}{dt}\] and x is the direction which the wire moves. it follows that speed is v. L is the length of the wire. B is the magnetic flux per unit area. It stays constant throughout. \[\frac{-d\Phi_{b}}{dt}=\frac{B\:dA}{dt}=\frac{BL\:dx}{dt}=BLv\]

OpenStudy (festinger):

I left out the minus sign, but it doesn't really matter in this case.

OpenStudy (anonymous):

wts dΦb? Flux density?

OpenStudy (festinger):

\[\Phi_{b}\] is magnetic flux. \[\frac{\Phi_{b}}{A}=B\] is magnetic flux density. \[\Phi_{b}=B*A\] B is magnetic field or magnetic flux density.

OpenStudy (anonymous):

ok. now i get the differentiation